OpenStudy (anonymous):

A new planet has been discovered, and it takes 84 years for it to orbit the Sun. What is the length of the semi-major axis of its orbit in AU? Enter the number, rounded to 1 decimal place and no units.

OpenStudy (anonymous):

Kepler's third law period=distance^(3/2) distance=period^(2/3) distance=82^(2/3) distance=18.9